OpenStudy (anonymous):

How to do 2 step equations

OpenStudy (anonymous):

do inverse operations of both steps example: 5x + 0 = 10 10 - 0 = 10 10/5 = 2 x = 2 is final answe. Also do the same to both sides

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcum